# Eigenvector decorrelation for random matrices

### v2

@@ 4.3.3. Conditional Gronwall estimates: regular case
$\begin{split}&\mathrm{Lin}^{(3)}_{t}=k(R_{1})\langle M_{12,t}^{R_{1}}\rangle(G_{1,t}G_{2,t}R_{2}G_{1,t}-M_{121,t}^{I,R_{2}})_{\bm{vw}}+k(R_{2})\langle M_{21,t}^{R_{2}}\rangle\big{(}G_{1,t}R_{1}G_{2,t}G_{1,t}-M_{121,t}^{R_{1},I}\big{)}_{\bm{vw}},\\[2.84526pt] &\mathrm{Err}^{(3)}_{1,t}=\big{\langle}G_{1,t}R_{1}G_{2,t}R_{2}G_{1,t}-M_{121,t}^{R_{1},R_{2}}\big{\rangle}\left(G_{1,t}^{2}\right)_{\bm{vw}}+\langle M_{121,t}^{R_{1},R_{2}}\rangle\left(G_{1,t}^{2}-M_{11,t}^{I}\right)_{\bm{vw}},\\[2.84526pt] &\mathrm{Err}^{(3)}_{2,t}=\big{\langle}G_{1,t}R_{1}G_{2,t}-M_{12,t}^{R_{1}}\big{\rangle}(G_{1,t}G_{2,t}R_{2}G_{1,t})_{\bm{vw}}+\big{\langle}G_{2,t}R_{2}G_{1,t}-M_{21,t}^{R_{2}}\big{\rangle}(G_{1,t}R_{1}G_{2,t}G_{1,t})_{\bm{vw}},\\[2.84526pt] &\mathrm{Err}^{(3)}_{3,t}=\langle G_{1,t}-M_{1,t}\rangle(G_{1,t}^{2}R_{1}G_{2,t}R_{2}G_{1,t})_{\bm{vw}}+\langle G_{2,t}-M_{2,t}\rangle(G_{1,t}R_{1}G_{2,t}^{2}R_{2}G_{1,t})_{\bm{vw}}\\[1.42262pt] &\qquad\qquad+\langle G_{1,t}-M_{1,t}\rangle(G_{1,t}R_{1}G_{2,t}R_{2}G_{1,t}^{2})_{\bm{vw}}\,.\end{split}$
@@
$\begin{split}&\tau^{R_{1}}:=\sup\left\{t\in[0,T]:\max_{s\in[0,t]}\max_{\bm{v},\bm{w}\in\{\bm{x},\bm{y}\}}\max_{z_{j,0}\in\Omega_{0}^{j}}\sqrt{N\ell_{s}\eta_{*,s}\gamma_{s}}\left|\left(G_{1,s}R_{1}G_{2,s}-M_{12,s}^{R_{1}}\right)_{\bm{vw}}\right|\leq N^{2\theta_{k(R_{1})}}\right\},\normalcolor\\ &\tau^{R_{1},R_{2}}:=\sup\left\{t\in[0,T]:\max_{s\in[0,t]}\max_{\bm{v},\bm{w}\in\{\bm{x},\bm{y}\}}\max_{z_{j,0}\in\Omega_{0}^{j}}\ell_{s}\gamma_{s}\left|\left(G_{1,s}R_{1}G_{2,s}R_{2}G_{1,s}^{(*)}\right)_{\bm{vw}}\right|\leq N^{2\xi_{k(R_{1},R_{2})}}\right\},\normalcolor\\ &\tau:=\min\left\{\tau^{R_{1}},\tau^{R_{1},R_{2}}:\,R_{1},R_{2}\in\mathfrak{S}\right\}\,,\quad\text{recalling}\quad\mathfrak{S}=\{I,B_{1},B_{1}^{*},B_{2},B_{2}^{*}\}\end{split}$
@@
(5.41a) $\left(\int_{0}^{t\wedge\tau}\mathrm{QV}_{s}^{(2)}\mathrm{d}s\right)^{1/2}+\int_{0}^{t\wedge\tau}|\mathcal{F}_{s}^{(2)}|\mathrm{d}s$ $\lesssim\frac{N^{\xi_{2k(R_{1})}}+k(R_{1})N^{2\theta_{0}}}{\sqrt{N\ell_{t\wedge\tau}\eta_{*,t\wedge\tau}\gamma_{t\wedge\tau}}}\log N,$
(5.41b) $\left(\int_{0}^{t\wedge\tau}\mathrm{QV}_{s}^{(3)}\mathrm{d}s\right)^{1/2}+\int_{0}^{t\wedge\tau}|\mathcal{F}_{s}^{(3)}|\mathrm{d}s$ $\lesssim\frac{\sum_{j=1,2}k(R_{j})N^{2\xi_{k(R_{3-j})}}+N^{\xi_{0}}+k(R_{1},R_{2})N^{\xi_{2}}}{\ell_{t\wedge\tau}\gamma_{t\wedge\tau}}\log N$
@@
(5.43) $\begin{split}\left(G_{1}^{*}R_{2}^{*}G_{2}^{*}R_{1}^{*}\Im G_{1}R_{1}G_{2}R_{2}G_{1}\right)_{\bm{ww}}&=\left(S\right)_{\bm{xx}}\\ &\leq N\langle S\rangle\lVert x\rVert^{2}=N\langle\Im G_{1}R_{1}|G_{2}|R_{1}^{*}\rangle(G_{1}^{*}R_{2}^{*}|G_{2}|R_{2}G_{1})_{\bm{ww}}\,.\end{split}$
@@
where (5.45a) corresponds to the propagation of the upper bound on the lhs. of (4.27b) and (5.45b) is the analog of (4.27c). The proof of (5.45a), (5.45b) is analogous to the proof of (5.18) and is based on splitting of the interval of integration into $[0,s_{*}]$ and $[s_{*},t\wedge\tau]$, where $s_{*}$ is defined in (5.20). The only difference is that in the regime $s\in[0,s_{*}]$ one needs to use the bound $\gamma_{s}\geq\eta_{j,s}/\rho_{j,s}$, $j\in[2]$.
@@
with the same notational convention about ()(*)( ∗ ) as in (5.46). It is easy to see that $\mathrm{Reg}^{(2)}$ satisfies the bound (5.47).
@@
(5.48) $\begin{split}\tau^{R_{1}}&:=\sup\left\{t\in[0,T]:\max_{s\in[0,t]}\max_{z_{j,0}\in\Omega^{j}_{\kappa,0}}\alpha_{1,s}^{-1}\left(\big{|}g_{s}^{\mathring{R}_{1}^{12},I}\big{|}+\big{|}g_{s}^{I,\mathring{R}_{1}^{21}}\big{|}\right)\leq N^{2\xi_{1}}\right\},\\ \tau^{R_{1},R_{2}}&:=\sup\left\{t\in[0,T]:\max_{s\in[0,t]}\max_{z_{j,0}\in\Omega^{j}_{\kappa,0}}\alpha_{2,s}^{-1}\big{|}g_{s}^{\mathring{R}_{1}^{12},\mathring{R}_{2}^{21}}\big{|}\leq N^{2\xi_{2}}\right\},\\ \tau&:=\min\left\{\tau^{R_{1}},\tau^{R_{1},R_{2}}:R_{1},R_{2}\in\{A_{1},A_{1}^{*},A_{2},A_{2}^{*}\}\right\}\end{split}$
for some small $100<\xi_{1}<\xi_{2}<\epsilon/$ 10, where
@@
The argument around (5.47) shows that whenever $\phi\neq 1$, it contributes with controllable and irrelevant error terms ${\rm Reg}^{(1)}_{t}$ and ${\rm Reg}^{(2)}_{t}$ to (5.46) and to the analogue of (5.46) in the case of two regularized observables, respectively. Hence, for simplicity we may assume that for initial conditions $\nu_{j,0}=(z_{j,0},D_{j,0})$, $j=1,2$, it holds that
@@
$\mathring{R}_{1}^{\nu_{1,t},\nu_{2,t}}=R_{1}^{\nu_{1,T},\nu_{2,T}}\quad\text{and}\quad\mathring{R}_{2}^{\nu_{2,t},\nu_{1,t}}=R_{2}^{\nu_{2,T},\nu_{1,T}}$
@@
(5.51) $\begin{split}\mathrm{QV}\big{[}g_{t}^{A_{1},I}\big{]}&\lesssim\frac{1}{N^{2}\eta_{1,t}}\langle\Im G_{1}A_{1}G_{2}\Im G_{1}A_{1}G_{2}^{*}\rangle\lesssim\frac{1}{N\eta_{1,t}^{2}}\langle\Im G_{1}A_{1}|G_{2}|A_{1}\rangle\langle\Im G_{1}|G_{2}|\rangle,\\ \mathrm{QV}\big{[}g_{t}^{A_{1},A_{2}}\big{]}&\lesssim\frac{1}{N^{2}\eta_{1,t}}\langle\Im G_{1}A_{1}G_{2}A_{2}\Im G_{1}A_{2}G_{2}^{*}A_{1}\rangle\lesssim\frac{1}{N\eta_{1,t}^{2}}\langle\Im G_{1}A_{1}|G_{2}|A_{1}\rangle\langle\Im G_{1}A_{2}|G_{2}|A_{2}\rangle,\end{split}$
where we used the reduction inequalities from (5.27). Here we restricted the argument to the case $R_{1}=A_{1}$ when one observable is regularized and $R_{1}=A_{1}$, $R_{2}=A_{2}$ when both are regularized. In general, one needs to consider $R_{1},R_{2}\in\{A_{1},A_{1}^{*},A_{2},A_{2}^{*}\}$, but all these cases are analogous to the one considered in (5.51) and thus omitted. Notice that in the rhs. of (5.51) also $|G|$ appeared. Products of traces with some $G$’s replaced by $|G|$ were already handled in (5.29) using the integral representation (5.28), however, the situation here is more delicate as we need to ensure that it is still possible to gain the additional smallness coming from $A_{1},A_{2}$ being regular along the whole vertical line (5.29). This analysis was already performed in full detail in [28, Eqs. (6.3)–(6.10)], we thus not repeat it here. We point out that in [28] this was done for fixed spectral parameters, however, given Lemma 3.3, the fact that $z_{t}$ now changes in time does not cause any complication as assuming (5.50) the notion of regularity does not depend on time. Proceeding as in [28, Eqs. (6.3)–(6.10)], using (3.10), (4.22a), and (4.27a), we thus conclude
@@
By the path-wise Burkholder-Davis-Gundy inequality (see [72, Appendix B.6, Eq. (18)] with $c=0$ for continuous martingale) we thus obtain
@@
In this section, we prove the Gronwall estimates from Section 4.3. Throughout their proofs, we will extensively use that, for a smooth function $f$ and $W_{t}$ solving (4.30), by Itô’s formula, it holds that
(6.1) $\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{E}f(W_{t})=-\frac{1}{2}\sum_{a,b}\mathbf{E}w_{ab}(r)(\partial_{ab}f)(W_{t})+\frac{1}{2}\sum_{a,b}\sum_{\alpha\in\{ab,ba\}}\kappa(ab,\alpha)\mathbf{E}(\partial_{ab}\partial_{\alpha}f)(W_{t})\,,$
and hence by a cumulant expansion (see, e.g., [51, Prop. 3.2])
(6.2) $\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{E}f(W_{t})=\sum_{k=2}^{K-1}\sum_{a,b}\sum_{\bm{\alpha}\in\{ab,ba\}^{k}}\frac{\kappa(ab,\bm{\alpha})}{k!}\mathbf{E}(\partial_{ab}\partial_{\bm{\alpha}}f)(W_{t})+\Omega_{K}$
with some explicit error term $\Omega_{K}$. Here, for a $k$-tuple of double indices $\bm{\alpha}=(\alpha_{1},...,\alpha_{k})$ we used the shorthand notation $\kappa(ab,(\alpha_{1},...,\alpha_{k}))=\kappa(w_{ab},w_{\alpha_{1}},...,w_{\alpha_{k}})$ for the joint cumulant of $w_{ab},w_{\alpha_{1}},...,w_{\alpha_{k}}$ and set $\partial_{\bm{\alpha}}=\partial_{w_{\alpha_{1}}}...\partial_{w_{\alpha_{k}}}$ and $\partial_{ab}=\partial_{w_{ab}}$.
In order to simplify the following presentation, we will henceforth assume that there is no difference for off-diagonal ($a\neq b$) and diagonal ($a=b$) cumulants $\kappa(ab,\bm{\alpha})$ in (6.2). The general case can be handled with straightforward minor modifications and is thus left to the reader.
@@
By Itô’s formula (6.1) for $f(W_{t})=|R_{t}|^{2p}$, a cumulant expansion (6.2) and using that $\kappa(ab,(\alpha_{1},...,\alpha_{k}))\lesssim N^{-(k+1)/2}$ in (6.2), we find that
(6.3) $\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{E}|R_{t}|^{2p}\lesssim\sum_{k=3}^{K}\frac{1}{N^{k/2}}\sum_{l=0}^{k}\left|\sum_{a,b}\mathbf{E}(\partial_{ab}^{l}\partial_{ba}^{k-l}|R_{t}|^{2p})\right|+\mathcal{O}(N^{-100p})\,.$
Here, we truncated the cumulant expansion at $K=\mathcal{O}(p)$ and used the trivial bound $\|G\|\leq\eta^{-1}$ to estimate the error term $\Omega_{K}$ from (6.2) in (6.3).
@@
(6.5) $N^{-3/2}\left|\sum_{a,b}M_{\bm{x}a}M_{bb}M_{aa}(M_{12}^{B_{1}})_{b\bm{y}}\right|\lesssim N^{-1/2}\sqrt{\sum_{a}|M_{\bm{x}a}|^{2}}\sqrt{\sum_{b}\big{|}(M_{12}^{B_{1}})_{b\bm{y}}\big{|}^{2}}\lesssim\frac{1}{\sqrt{N}\,\gamma}$
@@
$N^{-3/2}\left|\sum_{a,b}(G-M)_{\bm{x}a}M_{bb}M_{aa}(M_{12}^{B_{1}})_{b\bm{y}}\right|\prec N^{-1}\big{|}(G-M)_{\bm{x}\bm{m}}\big{|}\sqrt{\sum_{b}\big{|}(M_{12}^{B_{1}})_{b\bm{y}}\big{|}^{2}}\lesssim\frac{1}{N\sqrt{\eta}\,\gamma}\,,$
@@
$N^{-3/2}\left|\sum_{a,b}M_{\bm{x}a}(G-M)_{bb}M_{aa}(G_{1}B_{1}G_{2}-M_{12}^{B_{1}})_{b\bm{y}}\right|\prec\sqrt{\sum_{a}|M_{\bm{x}a}|^{2}}\frac{\mathcal{E}_{0}}{\sqrt{N\eta}}\lesssim\frac{\mathcal{E}_{0}}{\sqrt{N\eta}}\,.$
@@
$N^{-3/2}\left|\sum_{a,b}M_{\bm{x}a}(G-M)_{ab}(M_{12}^{B_{1}})_{bb}(G-M)_{a\bm{y}}\right|\prec\frac{1}{N\eta\gamma}\sqrt{\sum_{a}|M_{\bm{x}a}|^{2}}\lesssim\mathcal{E}_{1}$
@@
$N^{-3/2}\left|\sum_{a,b}(G-M)_{\bm{x}a}(G-M)_{ab}(G_{1}B_{1}G_{2}-M_{12}^{B_{1}})_{bb}(G-M)_{a\bm{y}}\right|\prec N^{1/2}\frac{\mathcal{E}_{0}}{(N\eta)^{3/2}}\lesssim\frac{\mathcal{E}_{1}}{\sqrt{N}\eta^{7/6}}\,,$
@@
(6.9) $\begin{split}N^{-3/2}&\left|\sum_{a,b}M_{\bm{x}a}(M_{12}^{B_{1}})_{b\bm{y}}(G-M)_{\bm{x}a}M_{bb}(G_{1}B_{1}G_{2}-M_{12}^{B_{1}})_{a\bm{y}}\right|\\ &\prec N^{-1}\sqrt{\sum_{a}|M_{\bm{x}a}|^{2}}\left|(M_{12}^{B_{1}})_{\bm{m}\bm{y}}\right|\frac{\mathcal{E}_{0}}{\sqrt{N\eta}}\lesssim N^{-1/2}\frac{\mathcal{E}_{0}}{\sqrt{N\eta}\gamma}\lesssim\mathcal{E}_{1}^{2}\,,\end{split}$
@@
$\begin{split}N^{-3/2}&\left|\sum_{a,b}(G-M)_{\bm{x}a}(G_{1}B_{1}G_{2}-M_{12}^{B_{1}})_{b\bm{y}}M_{\bm{x}b}(G_{1}B_{1}G_{2}-M_{12}^{B_{1}})_{aa}(G-M)_{a\bm{y}}\right|\\ &\prec\sqrt{\sum_{b}|M_{\bm{x}b}|^{2}}\,\frac{\mathcal{E}_{0}^{2}}{N\eta}\lesssim\frac{\mathcal{E}_{1}^{2}}{N\eta^{4/3}}\,.\end{split}$
@@
$\begin{split}N^{-3/2}&\left|\sum_{a,b}(G-M)_{\bm{x}a}(G_{1}B_{1}G_{2}-M_{12}^{B_{1}})_{b\bm{y}}M_{\bm{x}a}(M_{12}^{B_{1}})_{b\bm{y}}(G-M)_{\bm{x}b}(M_{12}^{B_{1}})_{a\bm{y}}\right|\\ &\prec N^{-1}\frac{\mathcal{E}_{0}}{N\eta\gamma^{2}}\lesssim\mathcal{E}_{1}^{3}\,.\end{split}$
@@
$N^{-3/2}\sum_{l=0}^{3}\left|\sum_{a,b}\mathbf{E}(\partial_{ab}^{l}\partial_{ba}^{k-l}|R|^{2p})\right|\lesssim\left(1+\frac{1}{\sqrt{N}\eta^{3/2}}\right)\left(|R|^{2p}+N^{\xi}\mathcal{E}_{1}^{2p}\right)\,.$

### v3

@@ 4.3.3. Conditional Gronwall estimates: regular case
$\begin{split}\mathrm{Lin}^{(3)}_{t}&=k(R_{1})\langle M_{12,t}^{R_{1}}\rangle(G_{1,t}G_{2,t}R_{2}G_{1,t}-M_{121,t}^{I,R_{2}})_{\bm{vw}}\\ &+k(R_{2})\langle M_{21,t}^{R_{2}}\rangle\big(G_{1,t}R_{1}G_{2,t}G_{1,t}-M_{121,t}^{R_{1},I}\big)_{\bm{vw}},\\[2.84526pt] \mathrm{Err}^{(3)}_{1,t}&=\big\langle G_{1,t}R_{1}G_{2,t}R_{2}G_{1,t}-M_{121,t}^{R_{1},R_{2}}\big\rangle\left(G_{1,t}^{2}\right)_{\bm{vw}}+\langle M_{121,t}^{R_{1},R_{2}}\rangle\left(G_{1,t}^{2}-M_{11,t}^{I}\right)_{\bm{vw}},\\[2.84526pt] \mathrm{Err}^{(3)}_{2,t}&=\big\langle G_{1,t}R_{1}G_{2,t}-M_{12,t}^{R_{1}}\big\rangle(G_{1,t}G_{2,t}R_{2}G_{1,t})_{\bm{vw}}\\ &+\big\langle G_{2,t}R_{2}G_{1,t}-M_{21,t}^{R_{2}}\big\rangle(G_{1,t}R_{1}G_{2,t}G_{1,t})_{\bm{vw}},\\[2.84526pt] \mathrm{Err}^{(3)}_{3,t}&=\langle G_{1,t}-M_{1,t}\rangle(G_{1,t}^{2}R_{1}G_{2,t}R_{2}G_{1,t})_{\bm{vw}}+\langle G_{2,t}-M_{2,t}\rangle(G_{1,t}R_{1}G_{2,t}^{2}R_{2}G_{1,t})_{\bm{vw}}\\[1.42262pt] &+\langle G_{1,t}-M_{1,t}\rangle(G_{1,t}R_{1}G_{2,t}R_{2}G_{1,t}^{2})_{\bm{vw}}\,.\end{split}$
@@
$\begin{split}\tau^{R_{1}}:=&\inf\Bigl\{t\in[0,T]:\\ &\max_{s\in[0,t]}\max_{\bm{v},\bm{w}\in\{\bm{x},\bm{y}\}}\max_{z_{j,0}\in\Omega_{0}^{j}}\sqrt{N\ell_{s}\eta_{*,s}\gamma_{s}}\left|\left(G_{1,s}R_{1}G_{2,s}-M_{12,s}^{R_{1}}\right)_{\bm{vw}}\right|\geq N^{2\theta_{k(R_{1})}}\Bigr\},\normalcolor\\ \tau^{R_{1},R_{2}}:=&\inf\Bigl\{t\in[0,T]:\\ &\max_{s\in[0,t]}\max_{\bm{v},\bm{w}\in\{\bm{x},\bm{y}\}}\max_{z_{j,0}\in\Omega_{0}^{j}}\ell_{s}\gamma_{s}\left|\left(G_{1,s}R_{1}G_{2,s}R_{2}G_{1,s}^{(*)}\right)_{\bm{vw}}\right|\geq N^{2\xi_{k(R_{1},R_{2})}}\Bigr\},\normalcolor\\ \tau:=&\min\left\{\tau^{R_{1}},\tau^{R_{1},R_{2}}:\,R_{1},R_{2}\in\mathfrak{S}\right\}\,,\quad\text{recalling}\quad\mathfrak{S}=\{I,B_{1},B_{1}^{*},B_{2},B_{2}^{*}\}\end{split}$
@@
(5.41a) $\left(\int_{0}^{t\wedge\tau}\hskip-5.69054pt\mathrm{QV}_{s}^{(2)}\mathrm{d}s\right)^{1/2}\hskip-11.38109pt+\hskip-2.84526pt\int_{0}^{t\wedge\tau}\hskip-5.69054pt|\mathcal{F}_{s}^{(2)}|\mathrm{d}s\lesssim\frac{N^{\xi_{2k(R_{1})}}+k(R_{1})N^{2\theta_{0}}}{\sqrt{N\ell_{t\wedge\tau}\eta_{*,t\wedge\tau}\gamma_{t\wedge\tau}}}\log N,$
(5.41b) $\left(\int_{0}^{t\wedge\tau}\hskip-5.69054pt\mathrm{QV}_{s}^{(3)}\mathrm{d}s\right)^{1/2}\hskip-11.38109pt+\hskip-2.84526pt\int_{0}^{t\wedge\tau}\hskip-5.69054pt|\mathcal{F}_{s}^{(3)}|\mathrm{d}s\lesssim\frac{\sum_{j=1,2}k(R_{j})N^{2\xi_{k(R_{3-j})}}\!+\!N^{\xi_{0}}\!+\!k(R_{1},R_{2})N^{\xi_{2}}}{\ell_{t\wedge\tau}\gamma_{t\wedge\tau}}\log N$
@@
(5.43) $\begin{split}&\left(G_{1}^{*}R_{2}^{*}G_{2}^{*}R_{1}^{*}\Im G_{1}R_{1}G_{2}R_{2}G_{1}\right)_{\bm{ww}}=\left(S\right)_{\bm{xx}}\\ &\quad\leq N\langle S\rangle\lVert x\rVert^{2}=N\langle\Im G_{1}R_{1}|G_{2}|R_{1}^{*}\rangle(G_{1}^{*}R_{2}^{*}|G_{2}|R_{2}G_{1})_{\bm{ww}}\,.\end{split}$
@@
where (5.45a) corresponds to the propagation of the upper bound on the lhs. of (4.27b) and (5.45b) is the analog of (4.27c). The proof of (5.45a), (5.45b) is analogous to the proof of (5.18) and is based on splitting of the interval of integration into $0,s_{*}]$ and $s_{*},t\wedge\tau]$, where $s_{*}$ is defined in (5.20). The only difference is that in the regime $s\in[0,s_{*}]$ one needs to use the bound $\gamma_{s}\geq\eta_{j,s}/\rho_{j,s}$, $j\in[2]$.
@@
with the same notational convention about ()(*) as in (5.46). It is easy to see that $\mathrm{Reg}^{(2)}$ satisfies the bound (5.47).
@@
(5.48) $\begin{split}\tau^{R_{1}}&:=\inf\left\{t\in[0,T]:\max_{s\in[0,t]}\max_{z_{j,0}\in\Omega^{j}_{\kappa,0}}\alpha_{1,s}^{-1}\left(\big|g_{s}^{\mathring{R}_{1}^{12},I}\big|+\big|g_{s}^{I,\mathring{R}_{1}^{21}}\big|\right)\geq N^{2\xi_{1}}\right\},\\ \tau^{R_{1},R_{2}}&:=\inf\left\{t\in[0,T]:\max_{s\in[0,t]}\max_{z_{j,0}\in\Omega^{j}_{\kappa,0}}\alpha_{2,s}^{-1}\big|g_{s}^{\mathring{R}_{1}^{12},\mathring{R}_{2}^{21}}\big|\geq N^{2\xi_{2}}\right\},\\ \tau&:=\min\left\{\tau^{R_{1}},\tau^{R_{1},R_{2}}:R_{1},R_{2}\in\{A_{1},A_{1}^{*},A_{2},A_{2}^{*}\}\right\}\end{split}$
for some small $0<\xi_{1}<\xi_{2}<\epsilon/10$, where
@@
The argument around (5.47) shows that whenever $\phi\neq 1$, it contributes with controllable and irrelevant error terms ${\mathrm{R}eg}^{(1)}_{t}$ and ${\mathrm{R}eg}^{(2)}_{t}$ to (5.46) and to the analogue of (5.46) in the case of two regularized observables, respectively. Hence, for simplicity we may assume that for initial conditions $\nu_{j,0}=(z_{j,0},D_{j,0})$, $j=1,2$, it holds that
@@
$\mathring{R}_{1}^{\nu_{1,t},\nu_{2,t}}=\mathring{R}_{1}^{\nu_{1,T},\nu_{2,T}}\quad\text{and}\quad\mathring{R}_{2}^{\nu_{2,t},\nu_{1,t}}=\mathring{R}_{2}^{\nu_{2,T},\nu_{1,T}}$
@@
(5.51) $\begin{split}\mathrm{QV}\big[g_{t}^{A_{1},I}\big]&\lesssim\frac{1}{N^{2}\eta_{1,t}^{2}}\langle\Im G_{1}A_{1}G_{2}\Im G_{1}A_{1}G_{2}^{*}\rangle\lesssim\frac{1}{N\eta_{1,t}^{2}}\langle\Im G_{1}A_{1}|G_{2}|A_{1}\rangle\langle\Im G_{1}|G_{2}|\rangle,\\ \mathrm{QV}\big[g_{t}^{A_{1},A_{2}}\big]&\lesssim\frac{1}{N^{2}\eta_{1,t}^{2}}\langle\Im G_{1}A_{1}G_{2}A_{2}\Im G_{1}A_{2}G_{2}^{*}A_{1}\rangle\\ &\lesssim\frac{1}{N\eta_{1,t}^{2}}\langle\Im G_{1}A_{1}|G_{2}|A_{1}\rangle\langle\Im G_{1}A_{2}|G_{2}|A_{2}\rangle,\end{split}$
where we used the reduction inequalities from (5.27). Here we restricted the argument to the case $R_{1}=A_{1}$ when one observable is regularized and $R_{1}=A_{1}$, $R_{2}=A_{2}$ when both are regularized. In general, one needs to consider $R_{1},R_{2}\in\{A_{1},A_{1}^{*},A_{2},A_{2}^{*}\}$, but all these cases are analogous to the one considered in (5.51) and thus omitted. Notice that in the rhs. of (5.51) also $|G|$ appeared. Products of traces with some $G$’s replaced by $|G|$ were already handled in (5.29) using the integral representation (5.28), however, the situation here is more delicate as we need to ensure that it is still possible to gain the additional smallness coming from $A_{1},A_{2}$ being regular along the whole vertical line (5.29). This analysis was already performed in full detail in [29, Eqs. (6.3)–(6.10)], we thus not repeat it here. We point out that in [29] this was done for fixed spectral parameters, however, given Lemma 3.3, the fact that $z_{t}$ now changes in time does not cause any complication as assuming (5.50) the notion of regularity does not depend on time. Proceeding as in [29, Eqs. (6.3)–(6.10)], using (3.10), (4.22a), and (4.27a), we thus conclude
@@
By the path-wise Burkholder-Davis-Gundy inequality (see [73, Appendix B.6, Eq. (18)] with $c=0$ for continuous martingale) we thus obtain
@@
In this section, we prove the Gronwall estimates from Section 4.3.
@@
By Ito’s formula and a cumulant expansion, we find that
(6.1) $\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{E}|R_{t}|^{2p}\lesssim\sum_{k=3}^{K}\frac{1}{N^{k/2}}\sum_{a,b}\sum_{l=0}^{k}\left|\mathbf{E}(\partial_{ab}^{l}\partial_{ba}^{k-l}|R_{t}|^{2p})\right|+\mathcal{O}(N^{-100p})$
where we truncated the expansion at $K=\mathcal{O}(p)$ and used the trivial bound $\|G\|\leq\eta^{-1}$ to estimate the error term in (6.1).
@@
(6.3) $N^{-3/2}\Big|\sum_{a,b}M_{\bm{x}a}M_{bb}M_{aa}(M_{12}^{B_{1}})_{b\bm{y}}\Big|\lesssim N^{-1/2}\sqrt{\sum_{a}|M_{\bm{x}a}|^{2}}\sqrt{\sum_{b}\big|(M_{12}^{B_{1}})_{b\bm{y}}\big|^{2}}\lesssim\frac{1}{\sqrt{N}\,\gamma}$
@@
$N^{-3/2}\!\Big|\sum_{a,b}(G-M)_{\bm{x}a}M_{bb}M_{aa}(M_{12}^{B_{1}})_{b\bm{y}}\Big|\!\!\prec N^{-1}\big|(G-M)_{\bm{x}\bm{m}}\big|\sqrt{\sum_{b}\big|(M_{12}^{B_{1}})_{b\bm{y}}\big|^{2}}\lesssim\frac{1}{N\sqrt{\eta}\,\gamma}\,$,
@@
$N^{-3/2}\Big|\sum_{a,b}M_{\bm{x}a}(G-M)_{bb}M_{aa}(G_{1}B_{1}G_{2}-M_{12}^{B_{1}})_{b\bm{y}}\Big|\prec\sqrt{\sum_{a}|M_{\bm{x}a}|^{2}}\frac{\mathcal{E}_{0}}{\sqrt{N\eta}}\lesssim\frac{\mathcal{E}_{0}}{\sqrt{N\eta}}\,.$
@@
$N^{-3/2}\Big|\sum_{a,b}M_{\bm{x}a}(G-M)_{ab}(M_{12}^{B_{1}})_{bb}(G-M)_{a\bm{y}}\Big|\prec\frac{1}{N\eta\gamma}\sqrt{\sum_{a}|M_{\bm{x}a}|^{2}}\lesssim\mathcal{E}_{1}$
@@
$N^{-3/2}\Big|\sum_{a,b}(G-M)_{\bm{x}a}(G-M)_{ab}(G_{1}B_{1}G_{2}-M_{12}^{B_{1}})_{bb}(G-M)_{a\bm{y}}\Big|\prec\frac{N^{1/2}\mathcal{E}_{0}}{(N\eta)^{3/2}}\lesssim\frac{\mathcal{E}_{1}}{\sqrt{N}\eta^{7/6}}\,$,
@@
[…]
@@
$\begin{split}N^{-3/2}&\Big|\sum_{a,b}(G-M)_{\bm{x}a}(G_{1}B_{1}G_{2}-M_{12}^{B_{1}})_{b\bm{y}}M_{\bm{x}b}(G_{1}B_{1}G_{2}-M_{12}^{B_{1}})_{aa}(G-M)_{a\bm{y}}\Big|\\ &\prec\sqrt{\sum_{b}|M_{\bm{x}b}|^{2}}\,\frac{\mathcal{E}_{0}^{2}}{N\eta}\lesssim\frac{\mathcal{E}_{1}^{2}}{N\eta^{4/3}}\,.\end{split}$
@@
$N^{-3/2}\Big|\!\sum_{a,b}(G-M)_{\bm{x}a}(G_{1}B_{1}G_{2}-M_{12}^{B_{1}})_{b\bm{y}}M_{\bm{x}a}(M_{12}^{B_{1}})_{b\bm{y}}(G-M)_{\bm{x}b}(M_{12}^{B_{1}})_{a\bm{y}}\Big|\!\!\prec\frac{N^{-1}\mathcal{E}_{0}}{N\eta\gamma^{2}}\!\lesssim\mathcal{E}_{1}^{3}\,.$
@@
$N^{-3/2}\sum_{a,b}\sum_{l=0}^{3}\left|\mathbf{E}(\partial_{ab}^{l}\partial_{ba}^{k-l}|R|^{2p})\right|\lesssim\left(1+\frac{1}{\sqrt{N}\eta^{3/2}}\right)\left(|R|^{2p}+N^{\xi}\mathcal{E}_{1}^{2p}\right)\,.$